Find the weighted estimate, p (with a line over it) to test the claim that p 1 = p 2. Use α = 0.05. Assume the samples are random and independent. Sample statistics: n 1 = 50, x 1 = 35, and n 2 = 60, x 2 = 40.
Answer #1

Given that,

From given values ,

= (x1 + x2) / (n1 + n2) = (35 + 40) / (50 + 60) = 75 / 110 = 0.682

the weighted estimate, p = 0.682